OpenStudy (redneckoutlaw):

Why are even and odd exponents significant when simplifying square root radical expressions?

OpenStudy (anonymous):

\[\sqrt{x ^{6}} \sqrt{x ^{5}}\]

OpenStudy (anonymous):

can't wait to see the answer to this

OpenStudy (anonymous):

can you simplify both of those

OpenStudy (redneckoutlaw):

Ok

OpenStudy (redneckoutlaw):

No only 6

OpenStudy (anonymous):

both can be simplified

OpenStudy (anonymous):

what is x^6 simplified

OpenStudy (redneckoutlaw):

Ummm 3

OpenStudy (anonymous):

\[x ^{3}\]

OpenStudy (anonymous):

now can you simplify x^5

OpenStudy (redneckoutlaw):

No?

OpenStudy (anonymous):

remember how jim said to simplify odd numbers?

OpenStudy (anonymous):

\[\frac{ 5 }{ 2 }\]

OpenStudy (anonymous):

\[2\frac{ 1 }{ 2 }\]

OpenStudy (redneckoutlaw):

Ok. but how you get the 2 under it?

OpenStudy (anonymous):

pair up the exponents to get them outside of the radical sign

OpenStudy (anonymous):

\[\frac{ 6 }{ 2 }\]

OpenStudy (anonymous):

\[x ^{3}\]

OpenStudy (redneckoutlaw):

Ok.

OpenStudy (anonymous):

\[x ^{2}\sqrt{x}\]

OpenStudy (zarkon):

\[\sqrt{x^6}\] does not simplify to \(x^3\)

OpenStudy (anonymous):

\[\sqrt{x ^{6}}=\sqrt{x ^{2}}×\sqrt{x ^{3}}\]

OpenStudy (redneckoutlaw):

0,o im confused....

OpenStudy (anonymous):

two pairs of 3 when simplifying exponents in radicals they go outside the radical in x^2=x

OpenStudy (redneckoutlaw):

So does x6 simplfy to 3?

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

plug -1 into \(\sqrt{x^6}\) and \(x^3\) what do you get?

OpenStudy (redneckoutlaw):

I dont know 0,0 my teacher didnt show me that way... And where do you get -1

OpenStudy (zarkon):

\[1=\sqrt{(-1)^6}\neq(-1)^3=-1\]

OpenStudy (zarkon):

therefore \[\sqrt{x^6}\neq x^3\]

OpenStudy (anonymous):

@Zarkon \[x ^{3}\sqrt{-1}\] is the same thing as \[\sqrt{-1^{6}}\]

OpenStudy (anonymous):

goes out as pairs of 2 in as multiples

OpenStudy (redneckoutlaw):

:?

OpenStudy (redneckoutlaw):

Holy hell im confused

OpenStudy (zarkon):

\[\Large\sqrt{x^6}=|x^3|\]

OpenStudy (anonymous):

@Zarkon sorry if im sounding rude right now but how would you simplify the radical x^6

OpenStudy (zarkon):

i did...it is \[\large|x^3|\]

OpenStudy (zarkon):

graph them on a calculator

OpenStudy (redneckoutlaw):

Im like confused as hell 0,0. Like ifyou took a kid, and spun him on his head for ten minutes with out stopping, put him on his feet and made a dog chase him.. Thats me

zepdrix (zepdrix):

XD

OpenStudy (redneckoutlaw):

Im a dead serious right now... Im the kid running around like hes drunk while look back at the dog about to crap my pants

OpenStudy (redneckoutlaw):

So if x6 isnt x3 then what is it?

zepdrix (zepdrix):

It ends up being \(\Large\rm x^3\) sometimes, and \(\Large\rm -x^3\) at other times. \(\Large\rm |x|=\sqrt{x^2}=\pm x\) When we apply an `even root` to an `even exponent`, we have to write it like this. Zarkon's example was good. Lemme see if I can come up with another one :o hmm

OpenStudy (redneckoutlaw):

So if it was a odd root, it would be the -x^3

zepdrix (zepdrix):

Ahh the website froze on me :( grr

OpenStudy (redneckoutlaw):

D:

zepdrix (zepdrix):

Naw, odd roots don't give us trouble :d They give us back the sign that we originally had. Example:\[\Large\rm \sqrt[3]{-8}=-2\]We started with a negative, and ended up with a negative. Good. The negative that you put in front of x^3 is telling us we should have the opposite sign of whatever we start with. no no no :o\[\Large\rm \sqrt[3]{x^{12}}=x^4\]No negative to worry about when taking an odd root.

zepdrix (zepdrix):

Oh oh I misread the question actually. They were asking about `even/odd powers` not `even/odd roots`. Hmm.

OpenStudy (redneckoutlaw):

So with the 3sqrt x^12, youdivide the 12 by the 3?

zepdrix (zepdrix):

Yes. We can rewrite the expression like this:\[\Large\rm \sqrt[3]{x^{12}}=x^{12/3}\]

zepdrix (zepdrix):

They specifically asked about `square` roots though. Which is only the second root. So this question is kinda strange...

OpenStudy (redneckoutlaw):

Ok, so techincally, this question hs to have a answer like this: If you had a odd power, your would have a answer in the exponent. But when you have a Even power, it will be one thing. for exaple. x^13 it would be \[x ^{5}\sqrt{x}\]

OpenStudy (redneckoutlaw):

Well my spelling and grammar and word choice is poor tonight. im super tired

zepdrix (zepdrix):

Hmm, yah I kinda like the way you're wording that. You'll always be left with something under the root when you're applying it to the odd powered expression. Odd power: \[\Large\rm \sqrt{x^{13}}=\sqrt{x^{12}}\cdot\sqrt{x}=\pm x^6\sqrt{x}\] Even power:\[\Large\rm \sqrt{x^{12}}=\pm x^6\]

OpenStudy (redneckoutlaw):

Yes, thats what i meant, sorry i just havnt slept in like 2 days, been stressing out cuz im moving, but yes i meant it like that. just didnt know a way of wording it.

zepdrix (zepdrix):

cool :)

OpenStudy (redneckoutlaw):

:D so pretty much what i have to explain to my teacher, is write something like your example down, and explain why even and odds, which something like my answer would be So the biggest part of why even and odd exponents are inportant are becuase with odd exponents you will be always left with something under the squareroot (More or less ouvariable) While even powers on the other hand, you will be left with nothing except the variable and the power

OpenStudy (redneckoutlaw):

A Power*

zepdrix (zepdrix):

Yehhhh, seems fine. You could even maybe comment on how umm.... Even powers, when you take a square root will always give you an `integer` exponent in your result. Example: \(\Large\rm \sqrt{x^6}=\pm x^3\) Odd powers will always leave you with a `fractional` exponent. Example: \(\Large\rm \sqrt{x^5}=x^{5/2}\) Just another idea >.< I'm not really sure what this is for, online assignment or something? Hard to say exactly what they're looking for. It's worded a little weird :)

OpenStudy (redneckoutlaw):

Ya thats what i said, do u mind helping with 2 more. Not sureif they are right

zepdrix (zepdrix):

Close this thread, it's getting too long. Open your question in a new thread. I'll stop by if I can c: Gotta do some stuff though.

OpenStudy (redneckoutlaw):

ok thamks